OpenStudy (anonymous):

Evaluate and simplify tan 15º.

OpenStudy (anonymous):

Do you just want answer

OpenStudy (anonymous):

tan(15) = tan(45 - 30) = [tan(45) - tan(30) ]/[ 1 + tan(30)tan(45) = (1 - 1/√3) /(1 + 1/√3) = (√3 - 1)/(√3 + 1) = (√3 - 1)^2 /(3 - 1) = 1/2 [3 + 1 - 2√3 ] = (2 - √3 )

OpenStudy (anonymous):

just split it up into parts thats all

OpenStudy (anonymous):

i really apreciate this thank you

OpenStudy (shubhamsrg):

for the alternate method, you may also use tan 30 = 2tan15 / (1- tan^2 15) a quadratic eqn in tan 15 which you should be able to solve easily.. but still the method above is shorter and sweeter! :)

OpenStudy (anonymous):

yes true man

OpenStudy (shubhamsrg):

or how about this :  tan 15 = x/ sqrt(3) now from angle bisector theorem, sqrt(3) /x = 2 /(1-x) you may solve for x here and thus solve for tan 15 .. there are more methods !! :P